# Asymptotic geometry of non-abelian Hodge theory and Riemann–Hilbert correspondence, rank three $\widetilde{E}_{6}$ case

### v1

@@ 1 Introduction and statement of the main result
The main characters of this paper are, on the one hand, the moduli space $\mathcal{M}_{\textrm{Dol}}(\alpha)$ of gauge-equivalence classes of rank $3$ Higgs bundles on the complex projective line $\mathbb{C}P^{1}$ with three logarithmic points (called the Dolbeault moduli space) and, on the other hand, the space $\mathcal{M}_{\textrm{B}}(\textbf{c})$ of representations of the fundamental group of the thrice-punctured $\mathbb{C}P^{1}$ in $\operatorname{SL}(3,\mathbb{C})$ up to overall conjugation (called the Betti moduli space or character variety). These two spaces are special instances corresponding to the extended root system $\widetilde{E}_{6}$ of a more general setup, where the genus of the underlying curve and the structure group of the objects may be arbitrary. It is known in general that the Dolbeault and Betti moduli spaces are diffeomorphic to each other via a composition of the non-abelian Hodge and Riemann–Hilbert correspondences. However, this diffeomorphism is highly transcendental, and extensive literature is devoted to understanding its properties, as well as to using it to translate known facts about one of these moduli spaces to new statements about the other one. For instance, it is known that these moduli spaces carry natural complex algebraic variety structures (of dimension $2$ in the $\widetilde{E}_{6}$ case), but the diffeomorphism between them is not compatible with their complex structures. Our goal is to investigate the asymptotic behavior of this diffeomorphism $\psi$ at infinity in the special case $\widetilde{E}_{6}$ specified above (i.e., genus 00, rank $3$, and with suitably chosen parameters to be specified later), in the hope that this particular case will shed light on some phenomena that continue to hold in the general case too. We may summarize our main result as:
@@
${\mathcal{M}_{\textrm{Dol}}\setminus h^{-1}(B_{R}(0))}$${\mathcal{M}_{B}\setminus\psi(h^{-1}(B_{R}(0)))}$${\mathcal{H}\setminus B_{R}(0)}$${\lvert\mathcal{D}\partial\mathcal{M}_{B}\rvert}$$\scriptstyle{\psi}$$\scriptstyle{H}$$\scriptstyle{S}$ (1)
@@
$\operatorname{NAHC}\colon\mathcal{M}_{\textrm{Dol}}(\alpha)\to\mathcal{M}_{\textrm{dR}}(\beta,\tau)$
and the Riemann Hilbert correspondence
$\operatorname{RH}\colon\mathcal{M}_{\textrm{dR}}(\beta,\tau)\to\mathcal{M}_{\textrm{B}}(\textbf{c})$
and $H$ is the Hitchin fibration over the Hitchin base $\mathcal{H}$. The Hitchin base will turn out to be a one-dimensional affine space in our case (see Section 2.3 for more details), and $\mathcal{H}\setminus B_{R}(0)$ is a neighbourhood of infinity in the base, where $R\gg 1$. Moreover $\lvert\mathcal{D}\partial\mathcal{M}_{B}\rvert$ denotes the body of the dual (or nerve) complex of the compactifying divisor of the Betti space, and $S$ is Simpson’s natural map from a neighborhood of infinity in $\mathcal{M}_{B}$ to the body of the nerve complex (see Section 2.1 and 6.2 for more details).
@@
In order to prove our result, we will employ the asymptotic abelianization approach used in [50] to establish the Geometric P=W conjecture in rank 2 corresponding to the Painlevé cases, suitably adapted to rank 3. An important technical difference with the rank $2$ case is that in [49][50] the second author made use of certain local models called fiducial solutions [11] that are only available in rank $2$. In the present article, however, we have found a way to get around finding specific model solutions around the branch points by making use of recent results of T. Mochizuki and the second author [37] for the analysis. This simplifies the presentation and clarifies the picture, thus paving the way for potential higher-rank and higher-dimensional generalizations of our viewpoint. As a consequence of the analysis, we find that the asymptotic behavior of $\psi$ depends on a decomposition into sectors of the Hitchin base, namely it is determined by different exponential terms in each sector around infinity. To be more precise, the absolute values of the trace coordinates are of exponential growth in the parameter, with a constant factor that are constant on a sector. Over all sectors, these constant factors are the integrals of the Liouville $1$-form along some loop on the spectral curve. This kind of behavior is an instance of the Stokes phenomenon, that we analyze in detail in this particular case. As for the phases of the trace coordinates, we find that they are asymptotically equal on each sector to the holonomy of the spectral sheaf along certain loops that we determine.
@@
This paper is organized as follows. In Section 2, we recall the background material necessary to explain our arguments. In Section 3, we first describe the $\operatorname{GL}(3,\mathbb{C})$ (which is equivalent in this case to $\operatorname{SL}(3,\mathbb{C})$) character variety of the three-punctured sphere in general, and prove Theorem 1.2. We then analyze trace coordinates on the character variety, going back to classical work of R. Fricke and F. Klein [13], and which were generalized for $\operatorname{GL}(3,\mathbb{C})$ character varieties by S. Lawton [28][29]. In Section 4, we use T. Mochizuki’s asymptotic abelianization technique [36] and equivariant method of T. Mochizuki and the second author [37] to give the large-scale analysis of harmonic bundles. In Section 5 we use T. Mochizuki’s solution of the Hitchin WKB problem on non-critical paths [36, Section 2.3] to describe the parallel transport matrices, and apply Riemann–Hilbert correspondence to the previous setup. Finally in Section 6, we investigate the asymptotic behaviour of the trace coordinates under the Riemann–Hilbert correspondence, analyze the emerging Stokes phenomenon, and prove Theorem 1.1.
@@ 2.1 Basic notations, definitions and results
A parabolic structure is a quasi-parabolic structure together with a choice of weight vectors $\underline{\alpha}_{P}$ at each $P$. It gives rise to an $\mathbb{R}$-filtration $\mathcal{P}_{*}$ of $\mathcal{E}$. We always assume that the Higgs field is weakly parabolic, meaning that $\theta:l_{P}^{i}\rightarrow l_{P}^{i}\otimes K(D)$ at each $P\in D$. The Higgs field is called strongly parabolic if $\theta:l_{P}^{i}\rightarrow l_{P}^{i+1}\otimes K(D)$ at each $P\in D$, i.e. the residue is nilpotent with respect to the filtration in that the action on the graded pieces $\textrm{gr}_{l}^{\boldsymbol{\cdot}}\mathcal{E}$ of $l_{P}^{\bullet}$ is trivial.
As usual, under stability of a Higgs bundle $(E,\theta,\overline{\partial}_{E})$ we mean that for any proper holomorphic subbundle $F\subset E$ which satisfies $\theta:F\rightarrow F\otimes K(D)$, the inequality $\mu(F)<\mu(E)$ holds, where $\mu(E)=\frac{\textrm{deg}E}{\textrm{rank}E}$ is the slope of the bundle ($\mu(F)$ defined similarly). In the parabolic setting, we speak about $\alpha$-stability, which depends on the weight vectors $\underline{\alpha}_{P}$, and means that for all $F$ satisfying the above conditions
$\frac{\textrm{pdeg}_{\underline{\alpha}}E}{\textrm{rank}E}>\frac{\textrm{pdeg}_{\underline{\alpha}}F}{\textrm{rank}F},$
@@
$\textrm{pdeg}_{\underline{\alpha}}E=\textrm{deg}E+\sum_{P\in D}\sum_{j=1}^{3}\alpha_{P}^{j}$ (3)
$\textrm{pdeg}_{\underline{\alpha}}F=\textrm{deg}F+\sum_{P\in D}\sum_{j=1}^{3}\alpha_{P}^{j}\cdot\textrm{dim}(({\left.\kern-1.2ptF\vphantom{\big{|}}\right|_{P}}\cap l_{P}^{j-1})/({\left.\kern-1.2ptF\vphantom{\big{|}}\right|_{P}}\cap l_{P}^{j}))$ (4)
@@
Let $\textrm{SL}(E)$ be the principal bundle of automorphisms of $E$ which induce the identity on $\operatorname{det}E$. Then the group of complex gauge transformations, denoted by $\mathcal{G}$, is the group of sections of $\textrm{SL}(E)$. Moreover its Lie algebra consists of the sections of $\mathfrak{sl}(E)$, the vector bundle of traceless endomorphisms of $E$. The gauge group $\mathcal{G}$ acts on the Higgs bundles via
@@
The moduli space of harmonic, $\alpha$-stable, strongly parabolic, meromorphic $\textrm{SL}(3,\mathbb{C})$-Higgs bundles, with at most logarithmic singularities, with given weight vectors $\underline{\alpha}_{P}$, up to the complex gauge action, is called the Dolbeault moduli space, denoted by $\mathcal{M}_{\textrm{Dol}}(\alpha)$.
@@
In case of logarithmic Higgs bundles, for $\theta$ and $h$ the so called tameness condition is satisfied, that is at each $P\in D$, $h$ admits a lift along any ray to $P$, which grows at most polynomially in the standard metric. (Here we consider $h$ as an equivariant harmonic map from the universal cover of the Riemann surface to the Hermitian symmetric space $\textrm{GL}(3,\mathbb{C})/\textrm{U}(3)$). For such a tame, harmonic bundle $(E,\theta,\overline{\partial}_{E},h)$ the connection
@@
is integrable, and $\nabla^{1,0}$ has regular singularities. Fix again for all $P\in D$ some $\underline{\beta}_{P}=(\beta_{P}^{1},\beta_{P}^{2},\beta_{P}^{3})$ parabolic weight vectors and the $\underline{\tau}_{P}=(\tau_{P}^{1},\tau_{P}^{2},\tau_{P}^{3})$ eigenvalues of the residue of the connection. The definition of $\beta$-stability and parabolic structure of the integrable connection is just the same as for the Higgs field (the parabolic structure of the underlying vector bundle is already given, see also [32]). We again require that (5) is compatible with the filtration, that is $(\textrm{res}_{P}\nabla-\tau_{P}^{j}\operatorname{id})(l_{P}^{j})\subset l_{P}^{j+1}$, for all $P\in D$ and $j=0,1,2$. If the eigenvalues $\tau_{P}^{j}$ are pairwise different that this implies that $\textrm{res}_{P}\nabla$ is diagonal with respect to some basis compatible with the filtration. The complex gauge group action on the space of connections is also inherited from the action on the space of Higgs bundles.
@@
The moduli space of $\beta$-stable, parabolic, integrable $\textrm{SL}(3,\mathbb{C})$-connections, with regular singularities at the punctures, with given weight vectors $\underline{\beta}_{P}$ and given residues $\underline{\tau}_{P}$ at each $P\in D$, up to the complex gauge action, is called the de Rham moduli space, denoted by $\mathcal{M}_{\textrm{dR}}(\beta,\tau)$.
@@
$\rho:\pi_{1}(\mathbb{C}P^{1}\setminus D,x_{0})\rightarrow\textrm{SL}(3,\mathbb{C})$
that is well-defined up to simultaneous conjugation by elements of $\textrm{PGL}(3,\mathbb{C})$ (corresponding to different choices of a basis of solutions at $x_{0}$). The eigenvalues, denoted by $\underline{c}_{P}=(c_{P}^{1},c_{P}^{2},c_{P}^{3})$, of the local monodromy around $P\in D$ are determined by $(\underline{\beta}_{P},\underline{\tau}_{P})$. (As a matter of fact, the local system admits a filtration and corresponding weights too, but we will not need this extra structure here.)
@@
The moduli space of the above described representations is called the Betti moduli space or character variety, denoted by $\mathcal{M}_{\textrm{B}}(\textbf{c})$.
@@
[…]
@@
    [3] The spaces $\mathcal{M}_{\textrm{Dol}}(\alpha)$, $\mathcal{M}_{\textrm{dR}}(\beta,\tau)$ and $\mathcal{M}_{\textrm{B}}(\textbf{c})$ are $\mathbb{C}$-analytic manifolds, and there exists a diffeomorphism
    $\operatorname{NAHC}:\mathcal{M}_{\textrm{Dol}}(\alpha)\rightarrow\mathcal{M}_{\textrm{dR}}(\beta,\tau)$
@@
    $\operatorname{RH}:\mathcal{M}_{\textrm{dR}}(\beta,\tau)\rightarrow\mathcal{M}_{\textrm{B}}(\textbf{c}),$

### v2

@@ 1 Introduction and statement of the main result
The main characters of this paper are, on the one hand, the moduli space $\mathcal{M}_{\operatorname{Dol}}(\alpha)$ of gauge-equivalence classes of rank $3$ Higgs bundles on the complex projective line $\mathbb{C}P^{1}$ with three logarithmic points (called the Dolbeault moduli space) and, on the other hand, the space $\mathcal{M}_{\operatorname{B}}(\textbf{c})$ of representations of the fundamental group of the thrice-punctured $\mathbb{C}P^{1}$ in $\operatorname{SL}(3,\mathbb{C})$ up to overall conjugation (called the Betti moduli space or character variety). These two spaces are special instances corresponding to the extended root system $\widetilde{E}_{6}$ of a more general setup, where the genus of the underlying curve and the structure group of the objects may be arbitrary. It is known in general that the Dolbeault and Betti moduli spaces are diffeomorphic to each other via a composition of the non-abelian Hodge and Riemann–Hilbert correspondences. However, this diffeomorphism is highly transcendental, and extensive literature is devoted to understanding its properties, as well as to using it to translate known facts about one of these moduli spaces to new statements about the other one. For instance, it is known that these moduli spaces carry natural complex algebraic variety structures (of dimension $2$ in the $\widetilde{E}_{6}$ case), but the diffeomorphism between them is not compatible with their complex structures. Our goal is to investigate the asymptotic behavior of this diffeomorphism $\psi$ at infinity in the special case $\widetilde{E}_{6}$ specified above (i.e., genus 00, rank $3$, and with suitably chosen parameters to be specified later), in the hope that this particular case will shed light on some phenomena that continue to hold in the general case too. We may summarize our main result as:
@@
${\mathcal{M}_{\operatorname{Dol}}\setminus H^{-1}(B_{R}(0))}$${\mathcal{M}_{\operatorname{B}}\setminus\psi(H^{-1}(B_{R}(0)))}$${\mathcal{H}\setminus B_{R}(0)}$${\lvert\mathcal{D}\partial\mathcal{M}_{B}\rvert}$$\scriptstyle{\psi}$$\scriptstyle{H}$$\scriptstyle{S}$ (1)
@@
$\operatorname{NAHC}\colon\mathcal{M}_{\operatorname{Dol}}(\alpha)\to\mathcal{M}_{\operatorname{dR}}(\beta,\tau)$
and the Riemann–Hilbert correspondence
$\operatorname{RH}\colon\mathcal{M}_{\operatorname{dR}}(\beta,\tau)\to\mathcal{M}_{\operatorname{B}}(\textbf{c})$
and $H$ is the Hitchin fibration over the Hitchin base $\mathcal{H}$. The Hitchin base will turn out to be a one-dimensional affine space in our case (see Section 2.3 for more details), and $\mathcal{H}\setminus B_{R}(0)$ is a neighbourhood of infinity in the base, where $R\gg 1$. Moreover $\lvert\mathcal{D}\partial\mathcal{M}_{B}\rvert$ denotes the body of the dual (or nerve) complex of the compactifying divisor of the Betti space, and $S$ is Simpson’s natural map from a neighborhood of infinity in $\mathcal{M}_{B}$ to the body of the nerve complex (see Section 2.1 and 6.3 for more details).
@@
In order to prove our result, we will employ the asymptotic abelianization approach used in [50] to establish the Geometric P=W conjecture in rank 2 corresponding to the Painlevé cases, suitably adapted to rank 3. An important technical difference with the rank $2$ case is that in [49][50] the second author made use of certain local models called fiducial solutions [11] that are only available in rank $2$. In the present article, however, we have found a way to get around finding specific model solutions around the branch points by making use of recent results of T. Mochizuki and the second author [37] for the analysis. This simplifies the presentation and clarifies the picture, thus paving the way for potential higher-rank and higher-dimensional generalizations of our viewpoint. As a consequence of the analysis, we find that the asymptotic behavior of $\psi$ depends on a decomposition into sectors of the Hitchin base, namely it is determined by different exponential terms in each sector around infinity. To be more precise, the absolute values of the trace coordinates are of exponential growth in the parameter, with the expression of their constant factors depending on the sector. Over all sectors, these constant factors are the integrals of the Liouville $1$-form along some loop on the spectral curve. As for the phases of the trace coordinates, we find that they are asymptotically equal on each sector to the holonomy of the spectral sheaf along certain loops that we determine. This then gives an unconditional proof of Theorem 1.1. The behaviour of the trace coordinates is an instance of the Stokes phenomenon, that we analyze in detail in this particular case. Unfortunately, the analysis breaks down in the Stokes directions. However, assuming that the same asymptotic formulas are valid in these directions too (Assumption 2), we can fully understand the asymptotic behaviour of the diffeomorphism, leading to a proof of Theorem 1.1 subject to this condition.
@@
This paper is organized as follows. In Section 2, we recall the background material necessary to explain our arguments. In Section 3, we first describe the $\operatorname{GL}(3,\mathbb{C})$ (which is equivalent in this case to $\operatorname{SL}(3,\mathbb{C})$) character variety of the three-punctured sphere in general, and prove Theorem 1.2. We then analyze trace coordinates on the character variety, going back to classical work of R. Fricke and F. Klein [13], and which were generalized for $\operatorname{GL}(3,\mathbb{C})$ character varieties by S. Lawton [28][29]. In Section 4, we use T. Mochizuki’s asymptotic abelianization technique [36] and equivariant method of T. Mochizuki and the second author [37] to give the large-scale analysis of harmonic bundles. In Section 5 we use T. Mochizuki’s solution of the Hitchin WKB problem on non-critical paths [36, Section 2.3] to describe the parallel transport matrices, and apply Riemann–Hilbert correspondence to the previous setup. Finally in Section 6, we investigate the asymptotic behaviour of the trace coordinates under the Riemann–Hilbert correspondence, analyze the emerging Stokes phenomenon. We prove Theorem 1.1 in Section 6.3 without assuming any further condition, and in Section 6.4 relying on Assumption 2. We believe that the conditional proof of the result may be of independent interest, because it conjecturally explains the global behaviour of the diffeomorphism. This is the reason we include Section 6.4 too.
@@ 2.1 Basic notations, definitions and results
A parabolic structure is a quasi-parabolic structure together with a choice of weight vectors $\underline{\alpha}_{P}$ at each $P$. It gives rise to an $\mathbb{R}$-filtration $\mathcal{P}_{*}$ of $\mathcal{E}$. We always assume that the Higgs field is weakly parabolic, meaning that $\theta:l_{P}^{i}\rightarrow l_{P}^{i}\otimes K(D)$ at each $P\in D$. The Higgs field is called strongly parabolic if $\theta:l_{P}^{i}\rightarrow l_{P}^{i+1}\otimes K(D)$ at each $P\in D$, i.e. the residue is nilpotent with respect to the filtration in that the action on the graded pieces $\operatorname{gr}_{l}^{\boldsymbol{\cdot}}\mathcal{E}$ of $l_{P}^{\bullet}$ is trivial.
As usual, under stability of a Higgs bundle $(E,\theta,\overline{\partial}_{E})$ we mean that for any proper holomorphic subbundle $F\subset E$ which satisfies $\theta:F\rightarrow F\otimes K(D)$, the inequality $\mu(F)<\mu(E)$ holds, where $\mu(E)=\frac{\operatorname{deg}E}{\operatorname{rank}E}$ is the slope of the bundle ($\mu(F)$ defined similarly). In the parabolic setting, we speak about $\alpha$-stability, which depends on the weight vectors $\underline{\alpha}_{P}$, and means that for all $F$ satisfying the above conditions
$\frac{\operatorname{pdeg}_{\underline{\alpha}}E}{\operatorname{rank}E}>\frac{\operatorname{pdeg}_{\underline{\alpha}}F}{\operatorname{rank}F},$
@@
$\operatorname{pdeg}_{\underline{\alpha}}E=\operatorname{deg}E+\sum_{P\in D}\sum_{j=1}^{3}\alpha_{P}^{j}$ (3)
$\operatorname{pdeg}_{\underline{\alpha}}F=\operatorname{deg}F+\sum_{P\in D}\sum_{j=1}^{3}\alpha_{P}^{j}\cdot\operatorname{dim}(({\left.\kern-1.2ptF\vphantom{\big{|}}\right|_{P}}\cap l_{P}^{j-1})/({\left.\kern-1.2ptF\vphantom{\big{|}}\right|_{P}}\cap l_{P}^{j}))$ (4)
@@
Let $\operatorname{SL}(E)$ be the principal bundle of automorphisms of $E$ which induce the identity on $\operatorname{det}E$. Then the group of complex gauge transformations, denoted by $\mathcal{G}$, is the group of sections of $\operatorname{SL}(E)$. Moreover its Lie algebra consists of the sections of $\mathfrak{sl}(E)$, the vector bundle of traceless endomorphisms of $E$. The gauge group $\mathcal{G}$ acts on the Higgs bundles via
@@
The moduli space of harmonic, $\alpha$-stable, strongly parabolic, meromorphic $\operatorname{SL}(3,\mathbb{C})$-Higgs bundles, with at most logarithmic singularities, with given weight vectors $\underline{\alpha}_{P}$, up to the complex gauge action, is called the Dolbeault moduli space, denoted by $\mathcal{M}_{\operatorname{Dol}}(\alpha)$.
@@
In case of logarithmic Higgs bundles, for $\theta$ and $h$ the so called tameness condition is satisfied, that is at each $P\in D$, $h$ admits a lift along any ray to $P$, which grows at most polynomially in the standard metric. (Here we consider $h$ as an equivariant harmonic map from the universal cover of the Riemann surface to the Hermitian symmetric space $\operatorname{GL}(3,\mathbb{C})/\operatorname{U}(3)$). For such a tame, harmonic bundle $(E,\theta,\overline{\partial}_{E},h)$ the connection
@@
is integrable, and $\nabla^{1,0}$ has regular singularities. Fix again for all $P\in D$ some $\underline{\beta}_{P}=(\beta_{P}^{1},\beta_{P}^{2},\beta_{P}^{3})$ parabolic weight vectors and the $\underline{\tau}_{P}=(\tau_{P}^{1},\tau_{P}^{2},\tau_{P}^{3})$ eigenvalues of the residue of the connection. The definition of $\beta$-stability and parabolic structure of the integrable connection is just the same as for the Higgs field (the parabolic structure of the underlying vector bundle is already given, see also [32]). We again require that (5) is compatible with the filtration, that is $(\operatorname{res}_{P}\nabla-\tau_{P}^{j}\operatorname{id})(l_{P}^{j})\subset l_{P}^{j+1}$, for all $P\in D$ and $j=0,1,2$. If the eigenvalues $\tau_{P}^{j}$ are pairwise different that this implies that $\operatorname{res}_{P}\nabla$ is diagonal with respect to some basis compatible with the filtration. The complex gauge group action on the space of connections is also inherited from the action on the space of Higgs bundles.
@@
The moduli space of $\beta$-stable, parabolic, integrable $\operatorname{SL}(3,\mathbb{C})$-connections, with regular singularities at the punctures, with given weight vectors $\underline{\beta}_{P}$ and given residues $\underline{\tau}_{P}$ at each $P\in D$, up to the complex gauge action, is called the de Rham moduli space, denoted by $\mathcal{M}_{\operatorname{dR}}(\beta,\tau)$.
@@
$\rho:\pi_{1}(\mathbb{C}P^{1}\setminus D,x_{0})\rightarrow\operatorname{SL}(3,\mathbb{C})$
that is well-defined up to simultaneous conjugation by elements of $\operatorname{PGL}(3,\mathbb{C})$ (corresponding to different choices of a basis of solutions at $x_{0}$). The eigenvalues, denoted by $\underline{c}_{P}=(c_{P}^{1},c_{P}^{2},c_{P}^{3})$, of the local monodromy around $P\in D$ are determined by $(\underline{\beta}_{P},\underline{\tau}_{P})$. (As a matter of fact, the local system admits a filtration and corresponding weights too, but we will not need this extra structure here.)
@@
The moduli space of the above described representations is called the Betti moduli space or character variety, denoted by $\mathcal{M}_{\operatorname{B}}(\textbf{c})$.
@@
$\alpha_{P}^{i}=\beta_{P}^{i}=\tau_{P}^{i},\hskip 5.69046pt\operatorname{and}\hskip 5.69046ptc_{P}^{i}=e^{-2\pi\sqrt{-1}\alpha_{P}^{i}},\forall P\in D,\forall i\in\{1,2,3\}.$
@@
    [3] The spaces $\mathcal{M}_{\operatorname{Dol}}(\alpha)$, $\mathcal{M}_{\operatorname{dR}}(\beta,\tau)$ and $\mathcal{M}_{\operatorname{B}}(\textbf{c})$ are $\mathbb{C}$-analytic manifolds, and there exists a diffeomorphism
    $\operatorname{NAHC}:\mathcal{M}_{\operatorname{Dol}}(\alpha)\rightarrow\mathcal{M}_{\operatorname{dR}}(\beta,\tau)$
@@
    $\operatorname{RH}:\mathcal{M}_{\operatorname{dR}}(\beta,\tau)\rightarrow\mathcal{M}_{\operatorname{B}}(\textbf{c}),$